Find the measure of the supplement of 120 degrees.
Find the measure of the supplement of 60 degrees.
Find the measure of the complement of 3 degrees.

Answer by nyc_function(2741)

Supplement of 120 degrees = 180 - 120 or 60 degrees.
Supplement of 60 degrees = 180 - 60 or 120 degrees.
Complement of 60 degrees = 90 - 3 or 87 degrees.